OpenStudy (anonymous):

A playground is being designed where children can interact with their friends in certain combinations. If there is 1 child, there can be 0 interactions. If there are 2 children, there can be 3 interactions. If there are 3 children, there can be 12 interactions. If there are 4 children, there can be 39 interactions. Which recursive equation represents the pattern? A. an = an – 1 + 3(n – 1) B. an = an – 1 + 3(n – 1) C. an = an – 1 + (3n – 1) D. an = an – 1 + (n – 1)^3

OpenStudy (anonymous):

@phi can you help?

OpenStudy (anonymous):

I think its B

OpenStudy (anonymous):

or A but I'm not pisitive

OpenStudy (anonymous):

What's the difference between A and B?

OpenStudy (phi):

Choices A and B (as written) are the same formula. I assume there is a typo ?

OpenStudy (anonymous):

no I copied and pasted it.

OpenStudy (anonymous):

shoot I see what i did... A. an = an – 1 + 3^(n – 1) B. an = an – 1 + 3(n – 1) C. an = an – 1 + (3n – 1) D. an = an – 1 + (n – 1)3

OpenStudy (anonymous):

D. an = an – 1 + (n – 1)^3